# von Neumann Orbit Equivalence

### v2

@@ 1. Introduction
The study of orbit equivalence can be motivated also from an entirely different point of view, being a measurable counterpart to quasi-isometry of groups. Gromov [Gro93] introduced measure equivalence (ME) for countable discrete groups as a measurable analogue of quasi-isometry and since then this notion has proven to be an important tool in geometric group theory with connections to ergodic theory and operator algebras. Two infinite countable discrete groups $\Gamma$ and $\Lambda$ are measure equivalent if there is an infinite measure space $(\Omega,m)$ with commuting, measure-preserving actions $\Gamma{\,\curvearrowright\,}(\Omega,m)$ and $\Lambda{\,\curvearrowright\,}(\Omega,m)$, so that both the actions admit finite-measure fundamental domains $Y,X\in\Omega$, that is, $m(Y),m(X)<\infty$ and
@@
$[\Gamma:\Lambda]_{\Omega}:=\frac{m(X)}{m(Y)}.$
Notably, measure equivalence was used by Furman in [Fur99a, Fur99b] to prove strong rigidity results for lattices in higher rank simple Lie groups. ME relates back to OE because of the following fact, observed by Zimmer and Furman: for two discrete groups $\Gamma$ and $\Lambda$, admitting free OE actions is equivalent to having an ME-coupling of index 1. Moreover, for OE groups, an ME-coupling can be chosen so that the fundamental domains coincide [Fur99b, Theorem 3.3].
@@
Remark 1.4.
If $(\mathcal{M},\mathrm{Tr})$ is a von Neumann coupling between $\Gamma$ and $\Lambda$ with $\Gamma$ and $\Lambda$ fundamentals domains $q$ and $p$, respectively, then the index of such a coupling is defined as
$[\Gamma:\Lambda]_{\mathcal{M}}:=\frac{\mathrm{Tr}(p)}{\mathrm{Tr}(q)}$
We suspect that the notion of vNE with coupling index 1 should be equivalent to the notion of vNE with common fundamental domain. However, we are unable to prove it at this point and leave it as an open problem.
Green [Gre90], in her Ph.D. thesis, introduced graph products of groups, another important group theoretical construction. If $\mathcal{G}=(V,E)$ is a simple, non-oriented graph with vertex set $V$ and edge set $E$, then the graph product of a family, $\{\Gamma_{v}\}_{v\in V}$, of groups indexed by $V$ is obtained from the free product $*_{v\in V}\Gamma_{v}$ by adding commutator relations determined by the edge set $E$. Depending on the graph, free products and direct products are special cases of the graph product construction. Adapting the ideas of [Gab05], Horbez and Huang [HH22, Proposition 4.2] proved the stability of OE under taking graph products over finite simple graphs (see also [Dem22]). To further explore the study of graph products within the context of measured group theory, we would like to draw the reader’s attention to the article [EH24]. In the current article, we also prove the stability of vNOE under taking graph products.
@@
In attempting to prove the above theorems, if one tries to adapt the techniques from one of [Gab05, HH22, Dem22], an immediate problem is presented by the lack of “point perspective” in the theory of von Neumann (orbit) equivalence. The lack of any natural non-commutative analogue of the notion of OE/ME cocycles, or that of measured equivalence relation can be considered as a few problems presented by the lack of point perspective. This often leads one to consider genuinely new techniques and different alternatives (see e.g., [IPR24, Ish24, Bat23a, Bat23b]). To overcome this obstruction, we introduce the notion of von Neumann orbit equivalence for tracial von Neumann algebras that is “compatible” with vNOE of groups (see Definition 1.6 and Theorem 3.10), and prove the analogues of Theorems 1.3 and 1.5 at the level of tracial von Neumann algebras.
@@ Acknowledgements.
The authors would like to express deep gratitude to Ben Hayes for his valuable insights, continuous support, and constant encouragement throughout this project. This work has benefitted profoundly from discussions with Ben Hayes on numerous occasions and from him encouraging us to look into [BV]. The project began with the authors’ meeting at the “NCGOA Spring Institute 2023” held at the Vanderbilt University, and we thank the organizers of the conference for the opportunity. A.Wu would also like to acknowledge the support from the NSF CAREER award DMS ##\##214473. Finally, we would like to thank the anonymous referee for helpful suggestions and pointing out an argument to us that helped in simplifying the proof of Theorem 1.8 and extending it to the case of an arbitrary countably infinite graph product.
@@ 2.1. Tracial von Neumann algebras and the standard form
A tracial von Neumann algebra $A$ is endowed with a trace, i.e., a faithful, normal, unital linear functional $\tau:A\to\mathbb{C}$ such that $\tau(xy)=\tau(yx)$ for all $x,y\in A$. The trace $\tau$ induces an inner product on $A$ given by $\langle x,y\rangle=\tau(y^{*}x),x,y\in A$, and we let $L^{2}(A)$ denote the Hilbert space completion of $A$ with respect to this inner product. When we view and element $x\in A$ as a vector in $L^{2}(A)$, we denote it by $\hat{x}$. An element $x\in A$ defines a bounded linear operator on $L^{2}(A)$ given by $L_{x}(\hat{y})=\widehat{xy},y\in A$, and thus we have a representation of $A$ in $\mathcal{B}(L^{2}(A))$, called the standard representation. There is also a canonical anti-linear conjugation operator $J:L^{2}(A)\to L^{2}(A)$, defined by $J(\hat{x})=\widehat{x^{*}},x\in A$, and we have that $A^{\prime}=JAJ$, where $A^{\prime}$ is the commutant of $M$ inside $\mathcal{B}(L^{2}(A))$. By the tracial property of $\tau$, for $x\in A$, the operator $R_{x}$ on $L^{2}(A)$ defined by $R_{x}(\hat{y})=\widehat{yx},y\in A$ is bounded. If we let $\rho(A)=\{R_{x}:x\in A\}\subset\mathcal{B}(L^{2}(A))$, then $\rho(A)=A^{\prime}$.
@@ 2.2. Free product and amalgamated free product
Let $(A_{1},\tau_{1})$ and $(A_{2},\tau_{2})$ be tracial von Neumann algebras. The free product of $A_{1}$ and $A_{2}$ is the unique, up to isomorphism, tracial von Neumann algebra $(A,\tau)$ containing $A_{1}$ and $A_{2}$, and is such that $\tau|_{A_{i}}=\tau_{i},i=1,2$, $A$ is generated by $A_{1}\cup A_{2}$, and $A_{1},A_{2}$ are free inside $A$, i.e., $\tau(a_{1}a_{2}\cdots a_{k})=0$, whenever $a_{i}\in A_{n_{i}}$ with $n_{1}\neq n_{2}\neq\cdots\neq n_{k},n_{i}\in\{1,2\}$, and $\tau_{n_{i}}(a_{i})=0$ for all $1\leq i\leq k$. We will say that an element $a_{1}a_{2}\cdots a_{k}$ of the algebraic free product $A_{1}*_{\mathrm{alg}}A_{2}$ is an alternating centered word with respect to $\tau$ if $a_{i}\in A_{n_{i}}$ with $n_{1}\neq n_{2}\neq\cdots\neq n_{k},n_{i}\in\{1,2\}$, and $\tau_{n_{i}}(a_{i})=0$ for all $1\leq i\leq k$.
For $i=1,2$, let $(A_{i},\tau_{i})$ be finite von Neumann algebras, $Q\subset A_{i}$ be a common von Neumann subalgebra, and $E_{i}:A_{i}\to Q$ be faithful, normal conditional expectations. The amalgamated free product $(A,E)=(A_{1},E_{1})*_{Q}(A_{2},E_{2})$ is a pair of a von Neumann algebra $A$ generated by $A_{1}$ and $A_{2}$ and a faithful normal conditional expectation $E:A\to Q$ such that $A_{1}$ and $A_{2}$ are freely independent with respect to $E$: $E(a_{1}a_{2}\cdots a_{k})=0$ whenever $a_{n_{i}}\in A_{n_{i}}$ with $n_{i}\in\{1,2,\},E_{n_{i}}(a_{i})=0$ and $n_{1}\neq n_{2}\neq\cdots\neq n_{k}.$ An element $a_{1}a_{2}\cdots a_{k}\in A$ will be called an alternating centered word with respect to $E$ if $a_{n_{i}}\in A_{n_{i}}$ with $n_{i}\in\{1,2\},E_{n_{i}}(a_{i})=0$ and $n_{1}\neq n_{2}\neq\cdots\neq n_{k}.$
@@ 2.4. Modules over tracial von Neumann algebras
    an $A-B$-bimodule is a triple $(\mathcal{H},\pi_{A},\pi_{B})$ such that $(\mathcal{H},\pi_{A})$ is a left $A$-module, $(\mathcal{H},\pi_{B})$ is a right $B$-module, and the representations $\pi_{A}$ and $\pi_{B}$ commute. For $\xi\in\mathcal{H}$, $x\in A$, and $y\in B$, we will write $x\xi y$ instead of $\pi_{A}(x)\pi_{B}(y)\xi~{}(=\pi_{B}(y)\pi_{A}(x)\xi).$
@@
Let $(A,\tau_{A}),(B,\tau_{B})$ be tracial von Neumann algebras and let $\mathcal{H}$ be an $A-B$-bimodule. A vector $\xi\in\mathcal{H}$ is called
@@
Suppose $(Q,\tau_{Q})$ is a tracial von Neumann algebra and $\mathcal{H}$ is a left $Q$-module. If $\xi\in\mathcal{H}$ is a tracial vector, then the orthogonal projection $P:\mathcal{H}\to\overline{\operatorname{Span}(Q\xi)}$ lies in ${}_{Q}\mathcal{B}(\mathcal{H})$. Moreover, since $\xi$ is tracial, the operator $U:L^{2}Q\to\overline{\operatorname{Span}(Q\xi)}$ given by $U\hat{x}=x\xi,x\in Q$ is a unitary. Extending $U$ to an isometry from $L^{2}Q$ into $\mathcal{H}$ in an obvious way and applying Proposition 2.4 to $T=PU:L^{2}Q\to\mathcal{H}$ yields that $\mathrm{Tr}(P)=\tau(1)=1$.
@@ 3.1. Von Neumann orbit equivalence for tracial von Neumann algebras
    There exists a finite von Neumann algebra $(Q,\tau_{Q})$ and a pointed $A\,\overline{\otimes}\,Q-B$-bimodule $(\mathcal{H},\xi)$ such that $\xi\in\mathcal{H}$ is a cyclic and (bi-)tracial vector for both $A\,\overline{\otimes}\,Q$-module structure, and $Q-B$-bimodule structure. That is, for all $a\in A,b\in B$, and $x\in Q$,
@@
For $b\in B$, we let $R_{b}\in\mathcal{B}(L^{2}(A\overline{\otimes}Q))$ be the operator corresponding to right multiplication by $b$. Since $\mathcal{H}$ is an $A\overline{\otimes}Q-B$ bimodule, and $U$ is $A\,\overline{\otimes}\,Q$-linear, so, the right action of $B$ commutes with the left action of $A\overline{\otimes}Q$ on $L^{2}(A\,\overline{\otimes}\,Q)$, and hence $R_{b}\in(A\,\overline{\otimes}\,Q)^{\prime}\cap\mathcal{B}(L^{2}(A\,\overline{\otimes}\,Q))$ for every $b\in B$. Since the commutant of $A\overline{\otimes}Q$ acting on $L^{2}(A\overline{\otimes}Q)$ is $\rho(A\overline{\otimes}Q)$ we define $\phi\colon B\to A\overline{\otimes}Q$ as follows: for $b\in B$, $\phi(b)$ is the unique element in $A\,\overline{\otimes}\,Q$ such that $R_{b}=\rho(\phi(b))$. This is directly checked to be a *-homomorphism. Moreover, by definition of $\phi$, we have that $\eta\cdot b=\eta\phi(b),$ for every $\eta\in L^{2}(A\,\overline{\otimes}\,Q),b\in B$. Since $\xi$ is $Q$-$B$ bi-tracial, we have for all $b\in B,x\in Q$ that
$\tau_{Q}(x)\tau_{B}(b)=\langle x\xi b,\xi\rangle=\langle U^{*}(x)b,U^{*}(\hat{1})\rangle=\langle U(U^{*}(x)b),\hat{1}\rangle=\langle x\cdot b,\hat{1}\rangle=\langle x\phi(b),\hat{1}\rangle=\tau(x\phi(b)),$
@@
$\langle x\xi\cdot b,\xi\rangle=\langle x\phi(b),\hat{1}\rangle=\tau(x\phi(b))=\tau_{Q}(x\mathbb{E}_{Q}(\phi(b)))=\tau_{Q}(x)\tau_{B}(b).$
@@
As mentioned in the introduction, the above definition depends on the choice of the traces $\tau_{A}$ and $\tau_{B}$. When in a situation where the traces are fixed and there is no risk of confusion, we will often drop them, and simply say that $A$ and $B$ are vNOE and write $A\sim_{\mathrm{vNOE}}B$.
@@
If $(A,\tau_{A})$ is a tracial von Neumann algebra, then taking $Q=\mathbb{C},\mathcal{H}=L^{2}(A)$, and $\xi=\hat{1}\in L^{2}(A)$ in condition (1) of Theorem 3.1 shows that $A\sim_{\mathrm{vNOE}}A$. To see symmetry, let $(A,\tau_{A})$ and $(B,\tau_{B})$ be tracial von Neumann algebras satisfying condition (1) of Theorem 3.1, and let $\mathcal{H},(Q,\tau_{Q}),\text{ and }\xi\in\mathcal{H}$ be as in the condition. Note that we can view $\mathcal{H}$ as an $A-B\,\overline{\otimes}\,Q^{\mathrm{op}}$-bimodule. Consider the conjugate Hilbert space $\overline{\mathcal{H}}$, and the corresponding canonical $B\,\overline{\otimes}\,Q^{\mathrm{op}}-A$ bimodule structure on $\overline{\mathcal{H}}$. Then, it is straightforward to check that the triple (¯$(\overline{\mathcal{H}},Q^{\mathrm{op}},\bar{\xi})$ satisfies (1) of Theorem 3.1 and thus, $B\sim_{\mathrm{vNOE}}A$. To show transitivity, we will use condition (2) of Theorem 3.1. To this end, let $(A,\tau_{A}),(B,\tau_{B})$, and $(C,\tau_{C})$ be tracial von Neumann algebras. Let $Q_{1},Q_{2}$, and $\phi_{1}:B\to A\,\overline{\otimes}\,Q_{1},\phi_{2}:C\to B\,\overline{\otimes}\,Q_{2}$ be as in (2) of Theorem 3.1. Let $Q=Q_{1}\,\overline{\otimes}\,Q_{2}$ and let $\phi:C\to A\,\overline{\otimes}\,Q$ be given by
@@
where $\phi_{1}\otimes\mathrm{id}_{Q_{1}}:B\,\overline{\otimes}\,Q_{2}\to A\,\overline{\otimes}\,Q_{1}\,\overline{\otimes}\,Q_{2}$ is the natural extension of $\phi_{1}:B\to A\,\overline{\otimes}\,Q_{1}$. Let $\mathbb{E}_{Q_{1}}:A\,\overline{\otimes}\,Q_{1}\to Q_{1},\mathbb{E}_{Q_{2}}:B\,\overline{\otimes}\,Q_{2}\to Q_{2},\text{and }\mathbb{E}_{Q}:A\,\overline{\otimes}\,Q_{1}\,\overline{\otimes}\,Q_{2}\to Q_{1}\,\overline{\otimes}\,Q_{2}$ be normal conditional expectations. Consider the map $\mathbb{E}_{Q_{1}}\otimes\mathrm{id}_{Q_{2}}:A\,\overline{\otimes}\,Q_{1}\,\overline{\otimes}\,Q_{2}\to Q_{1}\,\overline{\otimes}\,Q_{2}$. Note that $\mathbb{E}_{Q}=\mathbb{E}_{Q_{1}}\otimes\mathrm{\operatorname{id}}_{Q_{2}}$. Therefore,
@@
[…]
@@
Since $\overline{\operatorname{Span}\{\phi_{1}(b)x_{1}:b\in B,x_{1}\in Q_{1}\}}^{\|\cdot\|_{2}}=L^{2}(A\,\overline{\otimes}\,Q_{1})\supseteq A\,\overline{\otimes}\,Q_{1}$, the following computation completes the proof.
@@
$\supseteq\overline{\operatorname{Span}\{(\phi_{1}\otimes\operatorname{id}_{Q_{2}})((1\otimes x_{2})(\phi_{2}(c))(x_{1}\otimes 1):c\in C,x_{1}\in Q_{1},x_{2}\in Q_{2}\}}^{\|\cdot\|_{2}}$
@@
Checking $\overline{\operatorname{Span}\{\phi(b)x:b\in B,x\in Q\}}=L^{2}(A\overline{\otimes}Q)$ might not be easy in general. However, the following lemma simplifies verifying it in certain examples.

### v3

@@ 1. Introduction
The study of orbit equivalence can also be motivated from an entirely different point of view, being a measurable counterpart to quasi-isometry of groups. Gromov [Gro93] introduced measure equivalence (ME) for countable discrete groups as a measurable analogue of quasi-isometry and since then, this notion has proven to be an important tool in geometric group theory with connections to ergodic theory and operator algebras. Two infinite countable discrete groups $\Gamma$ and $\Lambda$ are measure equivalent if there is a standard infinite measure space $(\Omega,m)$ with commuting, measure-preserving actions $\Gamma{\,\curvearrowright\,}(\Omega,m)$ and $\Lambda{\,\curvearrowright\,}(\Omega,m)$, so that both the actions admit finite-measure fundamental domains $Y,X\in\Omega$, that is, $m(Y),m(X)<\infty$ and
@@
$\Gamma:\Lambda]_{\Omega}:=\frac{m(X)}{m(Y)}.$
Notably, measure equivalence was used by Furman in [Fur99a, Fur99b] to prove strong rigidity results for lattices in higher rank simple Lie groups. ME relates back to OE because of the following fact, observed by Zimmer and Furman: for two discrete groups $\Gamma$ and $\Lambda$, admitting free OE actions is equivalent to having an ME-coupling so that the fundamental domains coincide (see, for example, [Gab05, PME5], [Fur11, Theorem 2.5]).
@@
Green [Gre90], in her Ph.D. thesis, introduced graph products of groups, another important group-theoretical construction. If $\mathcal{G}=(V,E)$ is a simple, non-oriented graph with vertex set $V$ and edge set $E$, then the graph product of a family, $\{\Gamma_{v}\}_{v\in V}$, of groups indexed by $V$ is obtained from the free product $_{v\in V}\Gamma_{v}$ by adding commutator relations determined by the edge set $E$. Depending on the graph, free products and direct products are special cases of the graph product construction. Adapting the ideas of [Gab05], Horbez and Huang [HH22, Proposition 4.2] proved the stability of OE under taking graph products over finite simple graphs (see also [Dem22]). To further explore the study of graph products within the context of measured group theory, we would like to draw the reader’s attention to the article [EH24]. In the current article, we also prove the stability of vNOE under taking graph products.
@@
In attempting to prove the above theorems, if one tries to adapt the techniques from one of [Gab05, HH22, Dem22], an immediate problem is presented by the lack of “point perspective” in the theory of von Neumann (orbit) equivalence. The lack of any natural non-commutative analogue of the notion of OE/ME cocycles, or that of measured equivalence relation can be considered as a few problems presented by the lack of point perspective. This often leads one to consider genuinely new techniques and different alternatives (see e.g., [IPR24, Ish24, Bat23a, Bat23b]). To overcome this obstruction, we introduce the notion of von Neumann orbit equivalence for tracial von Neumann algebras that is “compatible” with vNOE of groups (see Definition 1.5 and Theorem 3.11), and prove the analogues of Theorems 1.3 and 1.4 at the level of tracial von Neumann algebras.
@@ Acknowledgements.
The authors would like to express deep gratitude to Ben Hayes for his valuable insights, continuous support, and constant encouragement throughout this project. This work has benefited profoundly from discussions with Ben Hayes on numerous occasions and from him encouraging us to look into [BV]. The project began with the authors’ meeting at the “NCGOA Spring Institute 2023” held at the Vanderbilt University, and we thank the organizers of the conference for the opportunity. A.Wu would also like to acknowledge the support from the NSF CAREER award DMS #\#214473. Finally, we would like to thank the anonymous referee for helpful suggestions and pointing out an argument to us that helped in simplifying the proof of Theorem 1.7 and extending it to the case of an arbitrary countably infinite graph product.
@@ 2.1. Tracial von Neumann algebras and the standard form
A tracial von Neumann algebra $A$ is endowed with a trace, i.e., a faithful, normal, unital linear functional $\tau:A\to\mathbb{C}$ such that $\tau(xy)=\tau(yx)$ for all $x,y\in A$. The trace $\tau$ induces an inner product on $A$ given by $\langle x,y\rangle=\tau(y^{*}x),x,y\in A$, and we let $L^{2}(A)$ denote the Hilbert space completion of $A$ with respect to this inner product. When we view an element $x\in A$ as a vector in $L^{2}(A)$, we denote it by $\hat{x}$. An element $x\in A$ defines a bounded linear operator on $L^{2}(A)$ given by $L_{x}(\hat{y})=\widehat{xy},y\in A$, and thus we have a representation of $A$ in $\mathcal{B}(L^{2}(A))$, called the standard representation. There is also a canonical anti-linear conjugation operator $J:L^{2}(A)\to L^{2}(A)$, defined by $J(\hat{x})=\widehat{x^{*}},x\in A$, and we have that $A^{\prime}=JAJ$, where $A^{\prime}$ is the commutant of $M$ inside $\mathcal{B}(L^{2}(A))$. By the tracial property of $\tau$, for $x\in A$, the operator $R_{x}$ on $L^{2}(A)$ defined by $R_{x}(\hat{y})=\widehat{yx},y\in A$ is bounded. If we let $\rho(A)=\{R_{x}:x\in A\}\subset\mathcal{B}(L^{2}(A))$, then $\rho(A)=A^{\prime}$.
@@ 2.2. Free product and amalgamated free product
Let $(A_{1},\tau_{1})$ and $(A_{2},\tau_{2})$ be tracial von Neumann algebras. The free product of $A_{1}$ and $A_{2}$ is the unique, up to isomorphism, tracial von Neumann algebra $(A,\tau)$ containing $A_{1}$ and $A_{2}$, and is such that $\tau|_{A_{i}}=\tau_{i},i=1,2$, $A$ is generated by $A_{1}\cup A_{2}$, and $A_{1},A_{2}$ are free inside $A$, i.e., $\tau(a_{1}a_{2}\cdots a_{k})=0$, whenever $a_{i}\in A_{n_{i}}$ with $n_{i}\neq n_{i+1}$ for all $i\in\{1,\ldots,k-1\}$, $n_{i}\in\{1,2\}$ and $\tau_{n_{i}}(a_{i})=0$ for all $1\leq i\leq k$. We will say that an element $a_{1}a_{2}\cdots a_{k}$ of the algebraic free product $A_{1}*_{\mathrm{alg}}A_{2}$ is an alternating centered word with respect to $\tau$ if $a_{i}\in A_{n_{i}}$ with $n_{i}\neq n_{i+1}$ for all $i\in\{1,\ldots,k-1\}$, $n_{i}\in\{1,2\}$ and $\tau_{n_{i}}(a_{i})=0$ for all $1\leq i\leq k$.
For $i=1,2$, let $(A_{i},\tau_{i})$ be tracial von Neumann algebras, $Q\subset A_{i}$ be a common von Neumann subalgebra, and $E_{i}:A_{i}\to Q$ be faithful, normal conditional expectations. The amalgamated free product $(A,E)=(A_{1},E_{1})*_{Q}(A_{2},E_{2})$ is a pair of a von Neumann algebra $A$ generated by $A_{1}$ and $A_{2}$ and a faithful normal conditional expectation $E:A\to Q$ such that $A_{1}$ and $A_{2}$ are freely independent with respect to $E$: $E(a_{1}a_{2}\cdots a_{k})=0$ whenever $a_{n_{i}}\in A_{n_{i}}$ with $n_{i}\in\{1,2,\},E_{n_{i}}(a_{i})=0$ for all $1\leq i\leq k$, and $n_{i}\neq n_{i+1}$ for all $i\in\{1,\ldots,k-1\}$. An element $a_{1}a_{2}\cdots a_{k}\in A$ will be called an alternating centered word with respect to $E$ if $a_{n_{i}}\in A_{n_{i}}$ with $n_{i}\in\{1,2\},E_{n_{i}}(a_{i})=0$ for all $1\leq i\leq k$, and $n_{i}\neq n_{i+1}$ for all $i\in\{1,\ldots,k-1\}$.
@@ 2.4. Modules over tracial von Neumann algebras
    an ABA-B-bimodule is a triple $(\mathcal{H},\pi_{A},\pi_{B})$ such that $(\mathcal{H},\pi_{A})$ is a left $A$-module, $(\mathcal{H},\pi_{B})$ is a right $B$-module, and the representations $\pi_{A}$ and $\pi_{B}$ commute. For $\xi\in\mathcal{H}$, $x\in A$, and $y\in B$, we will write $x\xi y$ instead of $\pi_{A}(x)\pi_{B}(y)\xi~(=\pi_{B}(y)\pi_{A}(x)\xi).$
@@
Let $(A,\tau_{A}),(B,\tau_{B})$ be tracial von Neumann algebras and let $\mathcal{H}$ be an ABA-B-bimodule. A vector $\xi\in\mathcal{H}$ is called
@@
Suppose $(Q,\tau_{Q})$ is a tracial von Neumann algebra and $\mathcal{H}$ is a left $Q$-module. If $\xi\in\mathcal{H}$ is a tracial vector, then the orthogonal projection $P:\mathcal{H}\to\overline{\operatorname{Span}(Q\xi)}$ lies in $_{Q}\mathcal{B}(\mathcal{H})$. Moreover, since $\xi$ is tracial, the operator $U:L^{2}Q\to\overline{\operatorname{Span}(Q\xi)}$ given by $U\hat{x}=x\xi,x\in Q$ is a unitary. Extending $U$ to an isometry from $L^{2}Q$ into $\mathcal{H}$ in an obvious way and applying Proposition 2.4 to $T=PU:L^{2}Q\to\mathcal{H}$ yields that $\mathrm{Tr}(P)=\tau_{Q}(1)=1$.
@@ 3.1. Von Neumann orbit equivalence for tracial von Neumann algebras
    There exists a tracial von Neumann algebra $(Q,\tau_{Q})$ and a pointed $A\,\overline{\otimes}\,Q-B$-bimodule $(\mathcal{H},\xi)$ such that $\xi\in\mathcal{H}$ is a cyclic and (bi-)tracial vector for both $A\,\overline{\otimes}\,Q$-module structure, and QBQ-B-bimodule structure. That is, for all $a\in A,b\in B$, and $x\in Q$,
@@
For $b\in B$, we let $R_{b}\in\mathcal{B}(L^{2}(A\overline{\otimes}Q))$ be the operator corresponding to right multiplication by $b$. Since $\mathcal{H}$ is an $A\overline{\otimes}Q-B$ bimodule, and $U$ is $A\,\overline{\otimes}\,Q$-linear, so, the right action of $B$ commutes with the left action of $A\overline{\otimes}Q$ on $L^{2}(A\,\overline{\otimes}\,Q)$, and hence $R_{b}\in(A\,\overline{\otimes}\,Q)^{\prime}\cap\mathcal{B}(L^{2}(A\,\overline{\otimes}\,Q))$ for every $b\in B$. Since the commutant of $A\overline{\otimes}Q$ acting on $L^{2}(A\overline{\otimes}Q)$ is $\rho(A\overline{\otimes}Q)$ we define $\phi\colon B\to A\overline{\otimes}Q$ as follows: for $b\in B$, $\phi(b)$ is the unique element in $A\,\overline{\otimes}\,Q$ such that $R_{b}=\rho(\phi(b))$. This is directly checked to be a normal *-homomorphism. Moreover, by definition of $\phi$, we have that $\eta\cdot b=\eta\phi(b),$ for every $\eta\in L^{2}(A\,\overline{\otimes}\,Q),b\in B$. Since $\xi$ is $Q$-$B$ bi-tracial, we have for all $b\in B,x\in Q$ that
$\tau_{Q}(x)\tau_{B}(b)=\langle x\xi b,\xi\rangle=\langle U^{*}(\hat{x})b,U^{*}(\hat{1})\rangle=\langle\widehat{x\phi(b)},\hat{1}\rangle=\tau(x\phi(b)),$
@@
$\langle x\xi\cdot b,\xi\rangle=\langle\widehat{x\phi(b)},\hat{1}\rangle=\tau(x\phi(b))=\tau_{Q}(x\mathbb{E}_{Q}(\phi(b)))=\tau_{Q}(x)\tau_{B}(b).$
@@
As mentioned in the introduction, the above definition depends on the choice of the traces $\tau_{A}$ and $\tau_{B}$. When in a situation where the traces are fixed, and there is no risk of confusion, we will often drop them, and simply say that $A$ and $B$ are vNOE and write $A\sim_{\mathrm{vNOE}}B$.
@@
If $(A,\tau_{A})$ is a tracial von Neumann algebra, then taking $Q=\mathbb{C},\mathcal{H}=L^{2}(A)$, and $\xi=\hat{1}\in L^{2}(A)$ in condition (1) of Theorem 3.1 shows that $A\sim_{\mathrm{vNOE}}A$. To see symmetry, let $(A,\tau_{A})$ and $(B,\tau_{B})$ be tracial von Neumann algebras satisfying condition (1) of Theorem 3.1, and let $\mathcal{H},(Q,\tau_{Q}),\text{ and }\xi\in\mathcal{H}$ be as in the condition. Note that we can view $\mathcal{H}$ as an $A-B\,\overline{\otimes}\,Q^{\mathrm{op}}$-bimodule. Consider the conjugate Hilbert space $\overline{\mathcal{H}}$, and the corresponding canonical $B\,\overline{\otimes}\,Q^{\mathrm{op}}-A$ bimodule structure on $\overline{\mathcal{H}}$. Then, it is straightforward to check that the triple $(\overline{\mathcal{H}},Q^{\mathrm{op}},\bar{\xi})$ satisfies (1) of Theorem 3.1 and thus, $B\sim_{\mathrm{vNOE}}A$. To show transitivity, we will use condition (2) of Theorem 3.1. To this end, let $(A,\tau_{A}),(B,\tau_{B})$, and $(C,\tau_{C})$ be tracial von Neumann algebras. Let $Q_{1},Q_{2}$, and $\phi_{1}:B\to A\,\overline{\otimes}\,Q_{1},\phi_{2}:C\to B\,\overline{\otimes}\,Q_{2}$ be as in (2) of Theorem 3.1. Let $Q=Q_{1}\,\overline{\otimes}\,Q_{2}$ and let $\phi:C\to A\,\overline{\otimes}\,Q$ be given by
@@
[…]
@@
$B\,\overline{\otimes}\,Q_{2}$$A\,\overline{\otimes}\,Q_{1}\,\overline{\otimes}\,Q_{2}$$Q_{2}$$Q_{1}\,\overline{\otimes}\,Q_{2}$$\scriptstyle{\mathbb{E}_{Q_{2}}}$$\scriptstyle{\phi_{1}\otimes\mathrm{id}_{Q_{2}}}$$\scriptstyle{\mathbb{E}_{Q_{1}}\otimes\mathrm{\operatorname{id}}_{Q_{2}}}$$\scriptstyle{1\otimes\mathrm{id}_{Q_{2}}}$
@@
Moreover, since $\phi_{1}$ is unital and $\phi_{1}\otimes\operatorname{id}_{Q_{2}}$ is a homomorphism, we also have that
$(\phi_{1}\otimes\operatorname{id}_{Q_{2}})(\phi_{2}(c)(1\otimes x_{2}))=(\phi_{1}\otimes\operatorname{id}_{Q_{2}})(\phi_{2}(c))(1\otimes x_{2})\quad\text{ for all }c\in C,x_{2}\in Q_{2}.$
Finally, since $\overline{\operatorname{Span}\{\phi_{1}(b)x_{1}:b\in B,x_{1}\in Q_{1}\}}^{\|\cdot\|_{2}}=L^{2}(A\,\overline{\otimes}\,Q_{1})\supseteq A\,\overline{\otimes}\,Q_{1}$, the following computation completes the proof.
@@
$\supseteq\overline{\operatorname{Span}\{(\phi_{1}\otimes\operatorname{id}_{Q_{2}})((\phi_{2}(c)(1\otimes x_{2}))(x_{1}\otimes 1):c\in C,x_{1}\in Q_{1},x_{2}\in Q_{2}\}}^{\|\cdot\|_{2}}$
@@
Checking $\overline{\operatorname{Span}\{\phi(b)x:b\in B,x\in Q\}}^{\|\cdot\|_{2}}=L^{2}(A\overline{\otimes}Q)$ might not be easy in general. However, the following lemma simplifies verifying it in certain examples.